OpenStudy (anonymous):

Solve for x: −3(x + 3) = −3(x + 1) − 5 A. 6 B. −6 C. All real numbers D. No solution

OpenStudy (ash2326):

We have \[-3(x+3)=-3(x+1)-5\] Let's simplify this, would you do that?

OpenStudy (anonymous):

I don't know how to do this...

OpenStudy (ash2326):

OK, I'll simplify the left side, you try the right Let's open the bracket on the left side \[-3\times (x+3)=-3(x+1)-5\] \[-3\times x-3\times 3=-3(x+1)-5\] we get \[-3x-9=-3(x+1)-5\] could you try the right side?